# Testing the tomographic Fermi liquid hypothesis with high-order cyclotron resonance

### v1

@@ Abstract
An electronic excitation in a three-dimensional Fermi liquid has a lifetime inversely proportional to the squared energy above the Fermi surface, $\tau_{ee}\propto\delta\varepsilon^{-2}$ Landau (1957). This fact follows from strong phase-space restrictions for electron-electron (e-e) scattering events. It enables the introduction of weakly interacting quasiparticles in the system of strongly interacting electrons, forming the basis of Fermi liquid (FL) theory Abrikosov and Khalatnikov (1959). Deviations from this scaling of electron lifetime are in the focus of modern condensed matter physics. Once the e-e scattering becomes strong, the conventional FL theory is no more able to predict the thermodynamic and kinetic properties. Situations favouring the breakdown of FL include reduced dimensionality Ishii et al. (2003), small Fermi surfaces Abrikosov and Beneslavskii (1971), flat bands Kumar et al. (2021); Sayyad et al. (2020), but are not limited to the latter.
@@
The above unconventional examples implied stronger carrier scattering, as compared to the normal FL case. Recently, it was shown that electron scattering in 2d electron systems (2DES) may be weaker than in normal FL Ledwith et al. (2019a). More precisely, the odd angular harmonics of the distribution function $\delta f_{m}\propto e^{im\theta}$ ($\theta$ is the momentum direction) were shown to relax via e-e scattering in an anomalously slow fashion. The scaling of odd-$m$ scattering rates follows a power-law dependence $\gamma_{m=2k+1}\equiv\tau^{-1}_{ee,m}\propto T^{\alpha}$ with $42<\alpha<$ 4 Kryhin and Levitov (2023); Hofmann and Gran (2023). The even-$m$ harmonics relax in a normal fashion $\gamma_{m=2k}\propto T^{2}\ln(\varepsilon_{F}/T)$ Zheng and Sarma (1996), where $\varepsilon_{F}$ is the Fermi energy. The physics beyond such relaxation hierarchy comes from the large phase space for collisions with zero total momentum of incoming particles (head-on collisions) in two dimensions, as compared to three dimensions Gurzhi et al. (1995). Yet, head-on collisions do not alter the odd-$m$ harmonics due to their anti-symmetry with respect to momentum ${\bf p}$. Another type of collisions, the low-angle collisions, relax only the distributions with strong angular dependence, i.e. $m\gg 1$.
@@
This new regime of carrier transport in two dimensions, dubbed as ’tomographic Fermi liquid’ (TFL), is challenging to test experimentally. The primary difficulty lies in the fact that e-e collisions do not directly affect electron mobility and resistivity in uniform fields and in large samples. They effect only on electron viscosity, the kinetic coefficient revealing itself only in highly non-uniform fields or small samples Ledwith et al. (2019a); Kryhin et al. (2023); Buhmann and Molenkamp (2002). As a result, the check of TFL hypothesis requires measurement of electron viscosity as a function of tuning parameters (e.g. sample width $W$ Ledwith et al. (2019a), frequency $\omega$ Hofmann and Das Sarma (2022) or temperature $T$ Kryhin et al. (2023)) and comparison of the obtained functional dependence with normal and tomographic FL theories. First, the measurement of viscosity is already challenging due to unavoidable residual scattering by impurities and phonons. Second, the NFL and TFL theories predict quite close scaling exponents for viscosity, which hinders the differentiation of these two regimes.
@@
Here, we show that high-order cyclotron resonance (CR) in the 2D electron systems Mohr and Heitmann (1982); Lefebvre et al. (1998); Bangert et al. (1996); Chaplik and Heitmann (1985) carries unique and unambiguous information about relaxation times of angular harmonics of the distribution function. More precisely, the width of $m$-th cyclotron resonance $\Gamma_{m}$ and the relaxation rate of the $m$-th distribution function harmonic $\gamma_{m}$ are equal, $\gamma_{m}=\Gamma_{m}$. This relation holds for weak scattering, $\gamma_{m}\ll\omega_{c}$, $\omega_{c}=e|B|/m$ is the cyclotron frequency, and for relatively uniform ac fields, $qR_{c}\ll 1$, $R_{c}$ is the cyclotron radius. In highly non-uniform fields with $qR_{c}\sim 1$, a more general relation between $\Gamma_{m}$ and $\gamma_{m}$ emerges. The measurement of width for sequential cyclotron peaks enables testing whether even and odd distribution harmonics relax at different rates.
@@
Prior theoretical studies of CR linewidth based on diagrammatic techniques Ando (1976); Ting et al. (1977); Kallin and Halperin (1985) overlooked this relation. A possible reason is the necessary account for conservation laws upon e-e collisions, which amounts to quite a challenging evaluation of vertex functions. In the classical magnetic fields, solution of kinetic equation provides a more transparent approach to the problem. Prior kinetic theories of CR assumed either identical relaxation rates for all angular harmonics Chaplik and Heitmann (1985), possibly excluding the relaxation of zeroth and first harmonics to account for conservation laws Pellegrino et al. (2017); Kapralov and Svintsov (2022), or were limited to the hydrodynamic approximation Müller and Sachdev (2008); Cruise et al. (2024). Suppression of the second CR harmonic by e-e collisions was evidenced in Ref. Mönch et al. (2023). Very recently, non-trivial modifications of cyclotron absorption due to the non-dissipative FL interactions were predicted Afanasiev et al. (2023). Here, we shall take a full account of different harmonic relaxation rates in the high-order CR problem.
@@
The relation between $m$-th order CR linewidth and relaxation rate of $m$-th angular harmonic can be understood as follows. The electromagnetic field changes its direction $m$ times during the $m$-th order cyclotron resonance, causing an $m$-fold deformation of the Fermi surface, $\delta f(\theta_{p})\propto e^{im\theta_{p}}$, see Fig. 1. Of course, high-order cyclotron absorption can occur only in non-uniform ac fields. This non-uniformity adds a phase factor $e^{iqR_{c}\sin\theta_{p}}$ to the distribution function (see Pitaevskii and Lifshitz (2012); Kapralov and Svintsov (2022) and Supplementary section I). Still, this phase factor makes no effect on linewidth in the limit $qR_{c}\ll 1$.
@@
Above, we have introduced the relaxation rates for $m$-th angular harmonics of the distribution function $\gamma_{m}=1/\tau_{m}$. The lowest three rates are zeros due to conservation of particle number and momentum upon collisions, $\gamma_{0,\pm 1}=0$ Abrikosov and Khalatnikov (1959). According to the tomographic Fermi liquid hypothesis, further harmonics satisfy $\{\gamma_{2k}\equiv\gamma_{\rm even}\}\gg\{\gamma_{2k+1}\equiv\gamma_{\rm odd}\}$.
Further on, it would be convenient to present the kinetic equation in the operator form similar to that used in quantum mechanics Ledwith et al. (2019a, b). We introduce the ket-vector for the distribution function |$\ket{\chi}$ related to the angular harmonics |$\ket{m}$ as |$\ket{\chi}=\sum{\chi_{m}\ket{m}}$, and the ket-vector for the electric forces |$\ket{F}$
@@
The calculated frequency-dependent conductivity of the tomographic Fermi liquid (8), is shown in Fig. 2  [^1] Application of Eq. 8 requires the knowledge of eigenvectors |$\ket{s}$ and eigenvalues $\omega_{s}+i\Gamma_{s}$ for the dynamic matrix. For arbitrary scattering strength, they can be found numerically by retaining a large (but finite) number of angular harmonics $m_{\max}$. The computational complexity here scales linearly with $m_{\max}$ due to the tridiagonal structure of the matrix. In the weak-scattering limit, the perturbative results (9,10) apply.. The real part of the conductivity displays sharp resonances at $\omega=m\omega_{c}$, $m\geq 1$. The cyclotron harmonics become stronger with increasing the field non-uniformity, i.e. at larger $q$. The main CR broadens at larger $q$, which is a consequence of the viscous character of damping in FL. So far, all these observations are in agreement with previous studies of CR in 2DES with weak carrier-carrier collisions Kapralov and Svintsov (2022); Pellegrino et al. (2017); Alekseev (2018).
@@
In the presence of spatial dispersion, the eigenvectors of ${\hat{H}}$ are spread across various angular harmonics (atomic sites) according to Saitoh (1973):
@@
The result of perturbation theory (10) reproduces very well the numerically obtained linewidths $\Gamma_{s}$ up to quite a large $qR_{c}\sim 10$, as shown in Fig. 3 (b). One observes that odd linewidths, being initially small, start growing at $qR_{c}\sim 1$. This occurs due to the ’admixture’ of the even angular harmonics to the odd cyclotron resonance. Remarkably, the linewidths of the even CRs (say, 2nd and 4th) can drop down to the very small values. This occurs due to the oscillatory nature of the spectral weight $|J_{0}(qR_{c})|^{2}$, and happens at $qR_{c}$ equal to the zeros of Bessel function $J_{0}$. We note finally that an expression for the viscous damping of the principal CR Alekseev (2018); Pellegrino et al. (2017); Kapralov and Svintsov (2022) immediately follows from a more general equation (10).
@@
Figure 4: Measured graphene photoresistance (PR) as a function of magnetic field for different values of temperature. The PR is defined as a change in a longitudinal resistance upon illumination $\Delta R_{xx}$, normalized by radiation intensity $I$. Panel (a) shows the $B$-dependent PR in the vicinity of 2nd, 3rd and 4th CR harmmonics at different temperatures (marked by numbers), panel (b) shows the magnified view of PR at $T=2$ K in the vicinity of 3rd and 4th harmonics. Solid lines show the fit according to Eq. (12), dashed vertical lines label the theoretically anticipated position of the CR. Radiation frequency is $f=0.69$ THz; carrier density is 3.9×3.9$\times$3.9 ×1012 cm-2 (c) Extracted widths of the 2nd (X2) and 3rd (X3) cyclotron resonances at different temperature. Green and orange data points were obtained at a different carrier density $n_{s}=3.25\times 10^{12}$ cm-2 and $T=4$ K (raw data in Supplementary section IV Sup )
@@
The experimental test of the TFL hypothesis now amounts to the measurement of the 2d conductivity at a given wave vector and frequency $\sigma(q,\omega)$. Its real part governs the absorbed power $P_{\rm abs}=1/2\sum_{q}{\sigma^{\prime}(q,\omega)|E_{q\omega}|^{2}}$, where $E_{q\omega}$ is the spectral composition of the electric ac field in the 2DES plane. One method for measuring $P_{\rm abs}$ relies on transmission spectroscopy of grating-gated 2DES Chaplik and Heitmann (1985). Such data are available only for the 2DES with modulated lateral doping Mohr and Heitmann (1982); Lefebvre et al. (1998), which complicates their analysis.
@@
Another powerful technique for measuring the absorption spectra of low-dimensional systems relies on the photoresistance Kukushkin et al. (2006); Kvon et al. (2008), i.e. the change in the $dc$ resistance $\Delta R_{\rm xx}$ induced by radiation. It relies on the fact that $\Delta R_{\rm xx}$ is proportional to the absorbed electromagnetic power, and resonances in absorption at certain $\omega$ and $B$ should greatly enhance the measured value of $\Delta R_{\rm xx}$.
@@
We have performed the photoresistance measurements on a high-quality graphene sample in classically strong magnetic fields $B<1$ T. The CR and its harmonics were excited by coherent THz radiation generated by a continuous wave optically pumped molecular laser. Details on sample fabrication and experimental technique can be found in Supplementary sections II and III Sup . We intentionally fabricated narrow ($\sim 1~{}\mu$m in width) metallic Hall contacts to the sample partially embedding into the channel. This configuration provides highly non-uniform electromagnetic fields and is thus particularly suitable for observation of high-order CRs Alonso-Gonzalez et al. (2014).
@@
Figure 4 (a,b) shows the example of the measured photoresistance at $f=0.69$ THz and $n_{s}=3.9\times$1012 cm-2, where up to three CR overtones ($s=2...4$) are observed. We note that the PR signal at the main $s=1$ CR is very weak at that frequency and density, but re-appears at larger $f$ and $n_{s}$ (Supplementary sections IV Sup  [^2] A possible reason for CR weakening at large $n_{s}$ lies in radiative dampingMuravev et al. (2016), i.e. screening of the incident radiation by highly conductive 2DES.. At given $n_{s}$ controlled by the back gate, we confidently identify all spikes to CR or its harmonics according to $B=2\pi mf/|e|s$, where $s$ is an integer and $m=\hbar\sqrt{\pi n_{s}}/v_{F}$ for single-layer graphene. The anticipated resonance positions are marked by vertical dashed lines in Fig. 4 (a-b), and agree well with observed photoresistance peaks. This agreement enables further linewidth analysis.
@@
To reproduce the highly asymmetric shape of the absorption, one has to account for the plasmonic effects, i.e. screening of the incident field $E_{0}$ by the 2d electrons Muravev et al. (2016). They are taken into account by relating the total field $E_{q\omega}$ to the incident field $E_{0}$ via the dielectric function of 2DES $\epsilon(q,\omega)$, $E_{q\omega}=E_{0}/\epsilon(q,\omega)$. The dominant contribution to the anti-screening comes from the waves with nearly-zero group velocity, the so-called Bernstein modes Dai et al. (2010); Dorozhkin et al. (2021); Volkov and Zabolotnykh (2014). In such situation, the spectral dependence of absorbed power is suitably described by Bandurin et al. (2022)
@@
Small scattering rate of the third harmonic of distribution function, as compared to the second one, cannot be explained within impurity or electron-phonon scattering models. It is easy to show that both these models predict stronger relaxation with increasing the harmonic number $m$ (see Baker et al. (2024) and Supplementary Section IV Sup ). The TFL hypothesis is currently the only one capable of explaining the unusual relation $\gamma_{3}<\gamma_{2}$. Still, it is worth noting that the $T$-dependence of relaxation time extracted from the experiment is not a power-law one. At least, the relaxation rates $\gamma_{2,3}$ have a large $T$-independent background which can be attributed to the built-in disorder.

### v2

@@ Abstract
An electronic excitation in a three-dimensional Fermi liquid has a lifetime inversely proportional to the squared energy above the Fermi surface, $\tau_{ee}\propto\delta\varepsilon^{-2}$ [1]. This fact follows from strong phase-space restrictions for electron-electron (e-e) scattering events. It enables the introduction of weakly interacting quasiparticles in the system of strongly interacting electrons, forming the basis of Fermi liquid (FL) theory [2]. Deviations from this scaling of electron lifetime are in the focus of modern condensed matter physics. Once the e-e scattering becomes strong, the conventional FL theory is no more able to predict the thermodynamic and kinetic properties. Situations favouring the breakdown of FL include reduced dimensionality [3], small Fermi surfaces [4], flat bands [5, 6], but are not limited to the latter.
@@
The above unconventional examples implied stronger carrier scattering, as compared to the normal FL case. Recently, it was shown that electron scattering in 2d electron systems (2DES) may be weaker than in normal FL [7]. More precisely, the odd angular harmonics of the distribution function $\delta f_{m}\propto e^{im\theta}$ ($\theta$ is the momentum direction) were shown to relax via e-e scattering in an anomalously slow fashion. The scaling of odd-$m$ scattering rates follows a power-law dependence $\gamma_{m=2k+1}\equiv\tau^{-1}_{ee,m}\propto T^{\alpha}$ with $42<\alpha<$ 4 [8, 9]. The even-$m$ harmonics relax in a normal fashion $\gamma_{m=2k}\propto T^{2}\ln(\varepsilon_{F}/T)$ [10], where $\varepsilon_{F}$ is the Fermi energy. The physics beyond such relaxation hierarchy comes from the large phase space for collisions with zero total momentum of incoming particles (head-on collisions) in two dimensions, as compared to three dimensions [11]. Yet, head-on collisions do not alter the odd-$m$ harmonics due to their anti-symmetry with respect to momentum ${\bf p}$. Another type of collisions, the low-angle collisions, relax only the distributions with strong angular dependence, i.e. $m\gg 1$.
@@
This new regime of carrier transport in two dimensions, dubbed as ’tomographic Fermi liquid’ (TFL), is challenging to test experimentally. The primary difficulty lies in the fact that e-e collisions do not directly affect electron mobility and resistivity in uniform fields and in large samples. They effect only on electron viscosity, the kinetic coefficient revealing itself only in highly non-uniform fields or small samples [7, 12, 13]. As a result, the check of TFL hypothesis requires measurement of electron viscosity as a function of tuning parameters (e.g. sample width $W$ [7], frequency $\omega$ [14] or temperature $T$ [12]) and comparison of the obtained functional dependence with normal and tomographic FL theories. First, the measurement of viscosity is already challenging due to unavoidable residual scattering by impurities and phonons. Second, the NFL and TFL theories predict quite close scaling exponents for viscosity, which hinders the differentiation of these two regimes.
@@
Here, we show that high-order cyclotron resonance (CR) in the 2D electron systems [15, 16, 17, 18] carries unique and unambiguous information about relaxation times of angular harmonics of the distribution function. More precisely, the width of $m$-th cyclotron resonance $\Gamma_{m}$ and the relaxation rate of the $m$-th distribution function harmonic $\gamma_{m}$ are equal, $\gamma_{m}=\Gamma_{m}$. This relation holds for weak scattering, $\gamma_{m}\ll\omega_{c}$, $\omega_{c}=e|B|/m$ is the cyclotron frequency, and for relatively uniform ac fields, $qR_{c}\ll 1$, $R_{c}$ is the cyclotron radius. In highly non-uniform fields with $qR_{c}\sim 1$, a more general relation between $\Gamma_{m}$ and $\gamma_{m}$ emerges. The measurement of width for sequential cyclotron peaks enables testing whether even and odd distribution harmonics relax at different rates.
@@
Prior theoretical studies of CR linewidth based on diagrammatic techniques [19, 20, 21] overlooked this relation. A possible reason is the necessary account for conservation laws upon e-e collisions, which amounts to quite a challenging evaluation of vertex functions. In the classical magnetic fields, solution of kinetic equation provides a more transparent approach to the problem. Prior kinetic theories of CR assumed either identical relaxation rates for all angular harmonics [18], possibly excluding the relaxation of zeroth and first harmonics to account for conservation laws [22, 23], or were limited to the hydrodynamic approximation [24, 25]. Suppression of the second CR harmonic by e-e collisions was evidenced in Ref. [26]. Very recently, non-trivial modifications of cyclotron absorption due to the non-dissipative FL interactions were predicted [27]. Here, we shall take a full account of different harmonic relaxation rates in the high-order CR problem.
@@
The relation between $m$-th order CR linewidth and relaxation rate of $m$-th angular harmonic can be understood as follows. The electromagnetic field changes its direction $m$ times during the $m$-th order cyclotron resonance, causing an $m$-fold deformation of the Fermi surface, $\delta f(\theta_{p})\propto e^{im\theta_{p}}$, see Fig. 1. Of course, high-order cyclotron absorption can occur only in non-uniform ac fields. This non-uniformity adds a phase factor $e^{iqR_{c}\sin\theta_{p}}$ to the distribution function (see [28, 23] and Appendix A). Still, this phase factor makes no effect on linewidth in the limit $qR_{c}\ll 1$.
@@
Above, we have introduced the relaxation rates for $m$-th angular harmonics of the distribution function $\gamma_{m}=1/\tau_{m}$. The lowest three rates are zeros due to conservation of particle number and momentum upon collisions, $\gamma_{0,\pm 1}=0$ [2]. According to the tomographic Fermi liquid hypothesis, further harmonics satisfy $\{\gamma_{2k}\equiv\gamma_{\rm even}\}\gg\{\gamma_{2k+1}\equiv\gamma_{\rm odd}\}$.
Further on, it would be convenient to present the kinetic equation in the operator form similar to that used in quantum mechanics [7, 29]. We introduce the ket-vector for the distribution function |$\ket{\chi}$ related to the angular harmonics |$\ket{m}$ as |$\ket{\chi}=\sum{\chi_{m}\ket{m}}$, and the ket-vector for the electric forces |$\ket{F}$
@@
The calculated frequency-dependent conductivity of the tomographic Fermi liquid (8), is shown in Fig. 2  [^1] Application of Eq. 8 requires the knowledge of eigenvectors |$\ket{s}$ and eigenvalues $\omega_{s}+i\Gamma_{s}$ for the dynamic matrix. For arbitrary scattering strength, they can be found numerically by retaining a large (but finite) number of angular harmonics $m_{\max}$. The computational complexity here scales linearly with $m_{\max}$ due to the tridiagonal structure of the matrix. In the weak-scattering limit, the perturbative results (9,10) apply.. The real part of the conductivity displays sharp resonances at $\omega=m\omega_{c}$, $m\geq 1$. The cyclotron harmonics become stronger with increasing the field non-uniformity, i.e. at larger $q$. The main CR broadens at larger $q$, which is a consequence of the viscous character of damping in FL. So far, all these observations are in agreement with previous studies of CR in 2DES with weak carrier-carrier collisions [23, 22, 31].
@@
In the presence of spatial dispersion, the eigenvectors of ${\hat{H}}$ are spread across various angular harmonics (atomic sites) according to [32]:
@@
The result of perturbation theory (10) reproduces very well the numerically obtained linewidths $\Gamma_{s}$ up to quite a large $qR_{c}\sim 10$, as shown in Fig. 3 (b). One observes that odd linewidths, being initially small, start growing at $qR_{c}\sim 1$. This occurs due to the ’admixture’ of the even angular harmonics to the odd cyclotron resonance. Remarkably, the linewidths of the even CRs (say, 2nd and 4th) can drop down to the very small values. This occurs due to the oscillatory nature of the spectral weight $|J_{0}(qR_{c})|^{2}$, and happens at $qR_{c}$ equal to the zeros of Bessel function $J_{0}$. We note finally that an expression for the viscous damping of the principal CR [31, 22, 23] immediately follows from a more general equation (10).
@@
Figure 4: Measured graphene photoresistance (PR) as a function of magnetic field for different values of temperature. The PR is defined as a change in a longitudinal resistance upon illumination $\Delta R_{xx}$, normalized by radiation intensity $I$. Panel (a) shows the $B$-dependent PR in the vicinity of 2nd, 3rd and 4th CR harmmonics at different temperatures (marked by numbers), panel (b) shows the magnified view of PR at $T=2$ K in the vicinity of 3rd and 4th harmonics. Solid lines show the fit according to Eq. (12), dashed vertical lines label the theoretically anticipated position of the CR. Radiation frequency is $f=0.69$ THz; carrier density is 3.9×3.9$\times$3.9 ×1012 cm-2 (c) Extracted widths of the 2nd (X2) and 3rd (X3) cyclotron resonances at different temperature. Green and orange data points were obtained at a different carrier density $n_{s}=3.25\times 10^{12}$ cm-2 and $T=4$ K (raw data in Appendix D)
@@
The experimental test of the TFL hypothesis now amounts to the measurement of the 2d conductivity at a given wave vector and frequency $\sigma(q,\omega)$. Its real part governs the absorbed power $P_{\rm abs}=1/2\sum_{q}{\sigma^{\prime}(q,\omega)|E_{q\omega}|^{2}}$, where $E_{q\omega}$ is the spectral composition of the electric ac field in the 2DES plane. One method for measuring $P_{\rm abs}$ relies on transmission spectroscopy of grating-gated 2DES [18]. Such data are available only for the 2DES with modulated lateral doping [15, 16], which complicates their analysis.
@@
Another powerful technique for measuring the absorption spectra of low-dimensional systems relies on the photoresistance [33, 34], i.e. the change in the $dc$ resistance $\Delta R_{\rm xx}$ induced by radiation. It relies on the fact that $\Delta R_{\rm xx}$ is proportional to the absorbed electromagnetic power, and resonances in absorption at certain $\omega$ and $B$ should greatly enhance the measured value of $\Delta R_{\rm xx}$.
@@
We have performed the photoresistance measurements on a high-quality graphene sample in classically strong magnetic fields $B<1$ T. The CR and its harmonics were excited by coherent THz radiation generated by a continuous wave optically pumped molecular laser. Details on sample fabrication and experimental technique can be found in Appendices B and C. We intentionally fabricated narrow ($\sim 1~{}\mu$m in width) metallic Hall contacts to the sample partially embedding into the channel. This configuration provides highly non-uniform electromagnetic fields and is thus particularly suitable for observation of high-order CRs [35].
@@
Figure 4 (a,b) shows the example of the measured photoresistance at $f=0.69$ THz and $n_{s}=3.9\times$1012 cm-2, where up to three CR overtones ($s=2...4$) are observed. We note that the PR signal at the main $s=1$ CR is very weak at that frequency and density, but re-appears at larger $f$ and $n_{s}$ (Appendix D)  [^2] A possible reason for CR weakening at large $n_{s}$ lies in radiative damping[37], i.e. screening of the incident radiation by highly conductive 2DES.. At given $n_{s}$ controlled by the back gate, we confidently identify all spikes to CR or its harmonics according to $B=2\pi mf/|e|s$, where $s$ is an integer and $m=\hbar\sqrt{\pi n_{s}}/v_{F}$ for single-layer graphene. The anticipated resonance positions are marked by vertical dashed lines in Fig. 4 (a-b), and agree well with observed photoresistance peaks. This agreement enables further linewidth analysis.
@@
To reproduce the highly asymmetric shape of the absorption, one has to account for the plasmonic effects, i.e. screening of the incident field $E_{0}$ by the 2d electrons [37]. They are taken into account by relating the total field $E_{q\omega}$ to the incident field $E_{0}$ via the dielectric function of 2DES $\epsilon(q,\omega)$, $E_{q\omega}=E_{0}/\epsilon(q,\omega)$. The dominant contribution to the anti-screening comes from the waves with nearly-zero group velocity, the so-called Bernstein modes [38, 39, 40]. In such situation, the spectral dependence of absorbed power is suitably described by [41]
@@
Small scattering rate of the third harmonic of distribution function, as compared to the second one, cannot be explained within impurity or electron-phonon scattering models. It is easy to show that both these models predict stronger relaxation with increasing the harmonic number $m$ (see [42] and Appendix E). The TFL hypothesis is currently the only one capable of explaining the unusual relation $\gamma_{3}<\gamma_{2}$. Still, it is worth noting that the $T$-dependence of relaxation time extracted from the experiment is not a power-law one. At least, the relaxation rates $\gamma_{2,3}$ have a large $T$-independent background which can be attributed to the built-in disorder.